OpenStudy (anonymous):

the ratio of an objects weight on earth to its weight on Neptune is 7:5. how much would a person who weighs 150 pounds on earth weight on Neptune?

OpenStudy (goformit100):

Welcome to Open Study ........................ Sir/Ma'am @denice3548

OpenStudy (anonymous):

7/5 = 150/x cross multiply (7)(x) = (5)(150) 7x = 750 x = 750/7 x = 107 1/7 lbs